OpenStudy (anonymous):

Find a power series representation for (4+2x)/(1+x) I know you have to do the power series for 1/(1+x) and then times it by (4+2x) but I don't know where to go from there.

OpenStudy (nottim):

Argh. So difficult. show any work you did, to help someone else help you.

OpenStudy (anonymous):

Well it's just (4+2x)*power series ((-1)^n*(x)^n) but the answer choices I have are just one power series (so no breaking the 4 and 2x up unless you combine power series), and some answer choices are 4 + power series

OpenStudy (zarkon):

\[\frac{4+2x}{1+x}=2+2\cdot\frac{1}{1+x}\]

OpenStudy (anonymous):

Wow I never would have seen that. Thank you so much.

OpenStudy (zarkon):

happy to help.